OpenStudy (anonymous):

A+B=750. B is 66 more than A. what are the values of A and B?

OpenStudy (anonymous):

so b = a + 66 a + (a + 66) = 750 2a + 66 = 750 2a = 684 a = 342. (342) + b = 750 b = 750 - 342 b = 408

OpenStudy (stamp):

B is 66 more than A B is, B = 66 more than A, A + 66 B is , 66 more than A B = A + 66

OpenStudy (anonymous):

i should remove the anwser i am sorry i should let him do!

OpenStudy (anonymous):

a+b = 750 b = a + 66 Substitute a+66 for b in the first equation a + (a + 66) = 750 Subtract 66 from each side of equation a + a = 684 2a = 684 Divide each side by 2 a = 342 If a = 342, b has to equal 408 for a+b =

OpenStudy (anonymous):

thankyou so much :)

OpenStudy (anonymous):

midal for me ^^

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

yea i give you the answer with detail :) before the other

OpenStudy (anonymous):

:D x

OpenStudy (anonymous):

the answer is 750(:

OpenStudy (anonymous):

hhhehehe you deserve it more :)

OpenStudy (anonymous):

do you see my answer it clear last time i didn't say anything but this time you don't deserve

OpenStudy (anonymous):

just because you you put 750 you want midal lol you are funny :P

OpenStudy (anonymous):

first of all nader1 can i say spell MEDAL right and secondly the oxford dictionary is only £6.99

OpenStudy (anonymous):

lmao;p

OpenStudy (anonymous):

well that explains it all, those frog legs must have gone to your brain and suck in real good

OpenStudy (anonymous):

thank you i am not English educated and i help you i am french educated maybe you can respect the other ! because i am here to help you not to learn English from you !

OpenStudy (anonymous):

lmfaooooo stop ! both of u.

OpenStudy (anonymous):

listen i don't want help people who don't respect other i but gonna report for that report abuse thank you and good luck no need from you the medal give it to seiga

OpenStudy (anonymous):

well sorry, atleast i know what to get you for your birthday... an English dictionary

OpenStudy (anonymous):

stop FIGHTING!

OpenStudy (anonymous):

@amistre64 @Preetha @SWAG

OpenStudy (anonymous):

#yolo

OpenStudy (anonymous):

seiga i am not fighting he is insulting people and he is new here !

OpenStudy (anonymous):

@Preetha U WERE BOTH FIGHTING, AND ITS NOT A HE. AND U WERE ACTING LIKE A JERK, IM GUNNA TELL THEM TO SUSPEND U.

OpenStudy (anonymous):

actually i am a she and you should know because chloe is a french name :/

OpenStudy (anonymous):

maybe you can respect other?

OpenStudy (anonymous):

look right i know im a donk but i was just so stressed !! :O

OpenStudy (anonymous):

i have become one of your fans xxx

OpenStudy (anonymous):

maybe you will not understand i gonna talk the moderator it better like that

OpenStudy (anonymous):

go for it

OpenStudy (anonymous):

no need to put me your fan because you are already be blocked good luck

OpenStudy (preetha):

Ok, both of you: Chloe was rude/flippant. Chloe apologized. Nader, please accept her apology like a gentleman and lets move on. Seiga, thanks for stepping in. Are we all good?

OpenStudy (anonymous):

i accept no problems

OpenStudy (anonymous):

thank you Preetha and good luck every one and anytime for any help :)

OpenStudy (anonymous):

and i remove the block from her for anytime she need help

OpenStudy (anonymous):

thanks guys :)

OpenStudy (anonymous):

no problems anytime